OpenStudy (anonymous):

i need help with 4/5*3/20?

OpenStudy (anonymous):

One way multiply the numbers on the top 4*3 = 12 multiply the numbers on the bottom 5*20 = 100 This gives 12/100 The greatest common factor or the biggest number we can divide 12 and 20 is 4 so 12 divided by 4 = 3 and 100 divided by 4 is 25 This fraction reduces to 4/25 Does that work for you?

OpenStudy (anonymous):

I'm not sure where Luis got 2 1/5.

OpenStudy (anonymous):

A slight fix. One way multiply the numbers on the top 4*3 = 12 multiply the numbers on the bottom 5*20 = 100 This gives 12/100 The greatest common factor or the biggest number we can divide 12 and 100 is 4 so 12 divided by 4 = 3 and 100 divided by 4 is 25 This fraction reduces to 4/25 Does that work for you?

OpenStudy (anonymous):

my teacher ended up with 55/72 how did he get that

OpenStudy (anonymous):

your teacher was working a different problem. the answer of 12/100 reduces to 3/25, despite the efforts of Luis and Transend to inject typos.